# The numeraire e-variable

### 29 Feb 2024 v1

@@ Abstract
We consider testing a composite null hypothesis ${\mathcal{P}}$ against a point alternative ${\mathbb{Q}}$. This paper establishes a powerful and general result: under no conditions whatsoever on ${\mathcal{P}}$ or ${\mathbb{Q}}$, we show that there exists a special e-variable $X^{*}$ that we call the numeraire. It is strictly positive and for every ${\mathbb{P}}\in{\mathcal{P}}$, ${\mathbb{E}}_{\mathbb{P}}[X^{*}]\leq 1$ (the e-variable property), while for every other e-variable $X$, we have ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq 1$ (the numeraire property). In particular, this implies ${\mathbb{E}}_{\mathbb{Q}}[\log(X/X^{*})]\leq 0$ (log-optimality). $X^{*}$ also identifies a particular sub-probability measure ${\mathbb{P}}^{*}$ via the density $d{\mathbb{P}}^{*}/d{\mathbb{Q}}=1/X^{*}$. As a result, $X^{*}$ can be seen as a generalized likelihood ratio of ${\mathbb{Q}}$ against ${\mathcal{P}}$. We show that ${\mathbb{P}}^{*}$ coincides with the reverse information projection (RIPr) when additional assumptions are made that are required for the latter to exist. Thus ${\mathbb{P}}^{*}$ is a natural definition of the RIPr in the absence of any assumptions on ${\mathcal{P}}$ or ${\mathbb{Q}}$. In addition to the abstract theory, we provide several tools for finding the numeraire in concrete cases. We discuss several nonparametric examples where we can indeed identify the numeraire, despite not having a reference measure. We end with a more general optimality theory that goes beyond the ubiquitous logarithmic utility. We focus on certain power utilities, leading to reverse Rényi projections in place of the RIPr, which also always exists.
@@ 1 Introduction
Suppose we observe data from an unknown distribution on a measurable space $(\Omega,{\mathcal{F}})$. This paper is concerned with testing the composite null hypothesis $H_{0}$ that the data distribution belongs to a given set ${\mathcal{P}}$ of distributions on this space, against the point alternative $H_{1}$ that the data was generated by a given distribution ${\mathbb{Q}}$. Here ${\mathcal{P}}$ and ${\mathbb{Q}}$ are completely arbitrary, and our general theory will not require any further assumptions whatsoever.
We approach the testing problem by designing e-variables: a $[0,\infty]$-valued random variable $X$ is called an e-variable for ${\mathcal{P}}$ if ${\mathbb{E}}_{\mathbb{P}}[X]\leq 1$ for all ${\mathbb{P}}\in{\mathcal{P}}$. Specifically, we will look for ‘optimal’ e-variables, in senses to be formalized later, that tend to be large under ${\mathbb{Q}}$.
E-variables measure evidence against the null: the larger its realized value, the stronger the evidence. One can also use e-variables to make hard accept/reject decisions. In particular, Markov’s inequality implies that we can reject $H_{0}$ at level $\alpha$ if an e-variable $X$ exceeds $1/\alpha$, since ${\mathbb{P}}(X\geq 1/\alpha)\leq\alpha$ for each ${\mathbb{P}}\in{\mathcal{P}}$. This paper will not be concerned with such hard decisions, but just in the measure of evidence $X$ itself, which requires no prespecification of any level $\alpha$.
@@
Our main result establishes that under absolutely no restrictions or conditions on ${\mathcal{P}}$ or ${\mathbb{Q}}$, there exists a special ${\mathbb{Q}}$-almost surely unique and strictly positive e-variable $X^{*}$ that we call the numeraire e-variable, or just the numeraire. It is characterized by the property that ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$. This in turn implies two other interpretable properties ${\mathbb{E}}_{\mathbb{Q}}[X^{*}/X]\geq 1$ and ${\mathbb{E}}_{\mathbb{Q}}[\log(X/X^{*})]\leq 0$, the second of which is called log-optimality. In fact, we show that $X^{*}$ is the numeraire if and only if it is log-optimal in this sense. The terminology derives from mathematical finance, where the numeraire portfolio is a central object analogous to the numeraire e-variable.
@@
Although no assumption on ${\mathcal{P}}$ or ${\mathbb{Q}}$ are needed in general, the theory simplifies if the alternative ${\mathbb{Q}}$ does not assign positive probability to any event that is null under every ${\mathbb{P}}\in{\mathcal{P}}$. In this case we say that ${\mathbb{Q}}$ is absolutely continuous with respect to ${\mathcal{P}}$, written ${\mathbb{Q}}\ll{\mathcal{P}}$. This natural generalization of absolute continuity of individual measures goes back to Halmos and Savage (1949) and is satisfied in most statistically relevant settings, while being significantly weaker than requiring a common dominating reference measure.
@@
The condition ${\mathbb{Q}}\ll{\mathcal{P}}$ has several consequences: (a) the numeraire is ${\mathbb{Q}}$-almost surely finite; (b) every e-variable is ${\mathbb{Q}}$-almost surely finite; and (c) the set of all e-variables for ${\mathcal{P}}$ is bounded in probability under ${\mathbb{Q}}$, meaning that $\lim_{a\to\infty}\sup_{X}{\mathbb{Q}}(X>a)=0$ where the supremum extends over all e-variables for ${\mathcal{P}}$. In fact, we show that the above properties are equivalent in that they imply each other, and they also imply that ${\mathbb{Q}}\ll{\mathcal{P}}$.
@@
Next, we associate to the numeraire a special sub-probability measure ${\mathbb{P}}^{*}\ll{\mathbb{Q}}$ defined via its density $d{\mathbb{P}}^{*}/d{\mathbb{Q}}=1/X^{*}$. This measure belongs to the bipolar of ${\mathcal{P}}$, denoted by ${\mathcal{P}}^{\circ\circ}$. The bipolar can be interpreted as the effective null hypothesis: it is the largest family of sub-probability measures that has the same set of e-variables as ${\mathcal{P}}$. These are the distributions against which e-variables for ${\mathcal{P}}$ do not provide any evidence.
@@
We show that ${\mathbb{P}}^{*}$ can be characterized in terms of properties normally associated with the reverse information projection (RIPr). Earlier work on the RIPr involve assumptions related to absolute continuity and/or finiteness of the minimum relative entropy between ${\mathbb{Q}}$ and ${\mathcal{P}}$, and in such settings the above ${\mathbb{P}}^{*}$ coincides with the RIPr. This includes, in particular, work by Grünwald et al. (2024) and Lardy et al. (2023), which in turn build on results in Li (1999). This leads us argue that ${\mathbb{P}}^{*}$ is the natural definition of the RIPr in the absence of any assumptions on ${\mathcal{P}}$ or ${\mathbb{Q}}$.
@@
One of our main contributions is to significantly generalize the conditions required for a log-optimality theory of e-variables to be built, and for a RIPr to exist. Furthermore, in addition to the abstract theory we develop tools to aid in the computation of the numeraire and the RIPr in concrete examples. We establish several sufficient conditions to verify that an e-variable at hand is indeed the numeraire, and we provide several examples of nonparametric nulls ${\mathcal{P}}$ (such as bounded distributions, symmetric distributions and sub-Gaussian distributions) that illustrate how these sufficient conditions can be used. These examples are out of the reach of previous works due to their lack of a reference measure on which the aforementioned works centrally rely.
@@
We end the paper with a more general optimality theory that goes beyond the ubiquitous logarithmic utility, by studying other concave utilities and associated divergences. Without aiming for full generality, we focus on certain power utilities in place of the logarithm, which leads to reverse Rényi projections in place of the RIPr. These developments, again, take place under absolutely no conditions on ${\mathcal{P}}$ or ${\mathbb{Q}}$.
@@
Finally, we note that if ${\mathcal{P}}=\{{\mathbb{P}}\}$ is a singleton and ${\mathbb{Q}}\ll{\mathbb{P}}$, we obtain as a consequence that the numeraire is the likelihood ratio $d{\mathbb{Q}}/d{\mathbb{P}}$, which is well known to be log-optimal (Shafer, 2021). Thus, our work can be viewed as deriving certain generalized likelihood ratios for composite null hypotheses, and extending the corresponding optimality theory.
@@
Section 2 defines the numeraire, establishes existence and uniqueness under no assumptions, characterizes the absolute continuity property ${\mathbb{Q}}\ll{\mathcal{P}}$, and establishes various other useful facts such as log-optimality of the numeraire. Section 3 introduces the RIPr and characterizes it via various optimality properties dual to the properties and log-optimality of the numeraire. Section 4 establishes a number of tools for finding the numeraire and the RIPr in practice. These tools are then applied to specific examples in Section 5. Finally, Section 6 goes beyond log-optimality by considering certain power utilities and the associated reverse Rényi projections. Section 7 concludes.
@@
Given the measurable space $(\Omega,{\mathcal{F}})$, we let $F_{+}$, $M_{+}$, and $M_{1}$ denote the set of all $[0,\infty]$-valued measurable functions, nonnegative measures, and probability measures on $\Omega$, respectively. Given a null hypothesis ${\mathcal{P}}$, which is an arbitrary nonempty subset of $M_{1}$, we write ${\mathcal{E}}$ for the associated set of e-variables, i.e., the family of all $X\in F_{+}$ that satisfy ${\mathbb{E}}_{\mathbb{P}}[X]\leq 1$ for all ${\mathbb{P}}\in{\mathcal{P}}$. Infinite quantities play a key role in this paper. We generally rely on standard measure theoretic conventions such as $00\cdot x=$ 0 for $x\in[-\infty,\infty]$ and $x\cdot\infty=\infty$ for $x\in(0,\infty]$, as well as $\log(0)=-\infty$ and $\log(\infty)=\infty$. When evaluating ratios of $[0,\infty]$-valued random variables we often use the following convention, the last part of which is somewhat unusual, but simplifies notation:
@@
Clarifying explanations are given in potentially ambiguous cases.
@@ 2 The numeraire e-variable
A numeraire e-variable (or just numeraire for short) is a ${\mathbb{Q}}$-almost surely strictly positive e-variable $X^{*}$ such that ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq 1$ for every e-variable $X$.
@@
Numeraires are unique up to ${\mathbb{Q}}$-nullsets. Indeed, if $X^{*}_{1}$ and $X^{*}_{2}$ are numeraires, then the ratio $Y=X^{*}_{2}/X^{*}_{1}$ satisfies $1\leq 1/{\mathbb{E}}_{\mathbb{Q}}[Y]\leq{\mathbb{E}}_{\mathbb{Q}}[1/Y]\leq 1$ thanks to the numeraire property of $X^{*}_{1}$, Jensen’s inequality, and the numeraire property of $X^{*}_{2}$. Thus Jensen’s inequality holds with equality, so $Y$ is ${\mathbb{Q}}$-almost surely equal to a constant which must be one. It follows that $X^{*}_{1}$ and $X^{*}_{2}$ are ${\mathbb{Q}}$-almost surely equal. In view of this uniqueness, we often speak of ‘the’ numeraire.
@@
In the case of a simple null $-~{\mathcal{P}}=\{\widetilde{\mathbb{P}}\}$ with ${\mathbb{Q}}$ absolutely continuous with respect to $-~\widetilde{\mathbb{P}}$, the numeraire is just the likelihood ratio $-~X^{*}=d{\mathbb{Q}}/d\widetilde{\mathbb{P}}$. Indeed, $X^{*}$ is ${\mathbb{Q}}$-almost surely strictly positive, it is an e-variable, and for any other e-variable $X$ we have ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]={\mathbb{E}}_{\widetilde{\mathbb{P}}}[X]\leq 1$. In particular settings, such as when a reference measure exists and certain additional assumptions are satisfied, we will show that the numeraire is a likelihood ratio involving the ‘reverse information projection’ (Grünwald et al., 2024; Lardy et al., 2023). Even without these restrictions, $X^{*}$ can still be interpreted as a likelihood ratio. One of our goals is to establish this in full generality. This is the subject of Section 3.
@@
If $X$ is an e-variable for ${\mathcal{P}}^{\prime}$, it is also an e-variable for ${\mathcal{P}}$ and one has ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq 1$ by assumption. This proves the lemma. ∎
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\log\frac{X}{X^{*}}\right]\leq 0$ for every e-variable $X$, (2.1)
where the left-hand side may be $\infty$- ∞. This follows directly from Jensen’s inequality and the numeraire property. The converse is also true, and we record the equivalence in the following proposition. The proof shows that the numeraire property is really the first-order condition for log-optimality. Note also that a numeraire $X^{*}$ is the ${\mathbb{Q}}$-almost surely unique log-optimal e-variable in the sense of (2.1) even if ${\mathbb{E}}_{\mathbb{Q}}[\log X^{*}]$ happens to be infinite.
@@
Let $X^{*}$ be a ${\mathbb{Q}}$-almost surely strictly positive e-variable. Then $X^{*}$ is a numeraire if and only if it is log-optimal. In particular, a log-optimal e-variable is unique up to ${\mathbb{Q}}$-nullsets.
@@
The forward direction was argued above. To prove the converse we assume $X^{*}$ is log-optimal. For any e-variable $X$ and $t\in(0,1)$, $X(t)=(1-t)X^{*}+tX$ is an e-variable. Thus by log-optimality, ${\mathbb{E}}_{\mathbb{Q}}[t^{-1}\log(X(t)/X^{*})]\leq 0$. The expression inside the expectation equals $t^{-1}\log(1-t+tX/X^{*})$, which converges to $X/X^{*}-1$ as $t$ tends to zero and is bounded below by $t^{-1}\log(1-t)$, hence by $-2\log 2$ for $t\in(0,1/2)$. Fatou’s lemma thus yields ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}-1]\leq 0$, showing that $X^{*}$ is a numeraire. Finally, the uniqueness statement follows from the equivalence just proved together with uniqueness of the numeraire up to ${\mathbb{Q}}$-nullsets. ∎
The following theorem shows that a numeraire always exists without any assumptions whatsoever on ${\mathcal{P}}$ or ${\mathbb{Q}}$. The theorem is related to an important family of results in the mathematical finance literature, known collectively as the Fundamental Theorem of Asset Pricing, which are versions of the statement that a numeraire portfolio with finite wealth exists if and only if unbounded profits must come with unbounded downside risk. The following theorem contains the corresponding statement in the context of e-variables.
@@
    The numeraire is ${\mathbb{Q}}$-almost surely finite.
@@
    Every e-variable is ${\mathbb{Q}}$-almost surely finite.
@@
    ${\mathbb{Q}}$ is absolutely continuous with respect to ${\mathcal{P}}$.
@@
    The set of e-variables ${\mathcal{E}}$ is bounded in probability under ${\mathbb{Q}}$.
@@
The existence of a numeraire is Lemma 2.9 below, which relies on the preceding Lemmas 2.7 and 2.8. Here we focus on the equivalence of the four conditions. Clearly (ii) implies (i). To see that (i) implies (ii), let $X^{*}$ be a ${\mathbb{Q}}$-almost surely finite numeraire and note that the numeraire property implies that any e-variable $X$ must be ${\mathbb{Q}}$-almost surely finite too. Next, we claim that (ii) is equivalent to (iii). Suppose (iii) fails. Then there is an event $A$ with ${\mathbb{Q}}(A)>0$ and ${\mathbb{P}}(A)=0$ for all ${\mathbb{P}}\in{\mathcal{P}}$, in which case $X=\infty\bm{1}_{A}$ is an e-variable that is not ${\mathbb{Q}}$-almost surely finite. Thus (ii) fails also. The converse follows by noting that every e-variable is finite ${\mathbb{P}}$-almost surely for all ${\mathbb{P}}\in{\mathcal{P}}$. Further, (iv) implies (ii) because under the former condition we have for any e-variable $X$ that ${\mathbb{Q}}(X=\infty)=\lim_{a\to\infty}{\mathbb{Q}}(X>a)=0$. It remains to prove that (ii) implies (iv). This more involved implication is the content of Lemma 2.6 below. ∎
@@
The rest of this section contains the remaining parts of the proof of Theorem 2.5.
@@
If every e-variable is ${\mathbb{Q}}$-almost surely finite then ${\mathcal{E}}$ is bounded in probability under ${\mathbb{Q}}$.

### 22 Mar 2024 v2

@@ Abstract
We consider testing a composite null hypothesis ${\mathcal{P}}$ against a point alternative ${\mathsf{Q}}$. This paper establishes a powerful and general result: under no conditions whatsoever on ${\mathcal{P}}$ or ${\mathsf{Q}}$, there exists a special e-variable $X^{*}$ that we call the numeraire. It is strictly positive and for every ${\mathsf{P}}\in{\mathcal{P}}$, ${\mathbb{E}}_{\mathsf{P}}[X^{*}]\leq 1$ (the e-variable property), while for every other e-variable $X$, we have ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ (the numeraire property). In particular, this implies ${\mathbb{E}}_{\mathsf{Q}}[\log(X/X^{*})]\leq 0$ (log-optimality). $X^{*}$ also identifies a particular sub-probability measure ${\mathsf{P}}^{*}$ via the density $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. As a result, $X^{*}$ can be seen as a generalized likelihood ratio of ${\mathsf{Q}}$ against ${\mathcal{P}}$. We show that ${\mathsf{P}}^{*}$ coincides with the reverse information projection (RIPr) when additional assumptions are made that are required for the latter to exist. Thus ${\mathsf{P}}^{*}$ is a natural definition of the RIPr in the absence of any assumptions on ${\mathcal{P}}$ or ${\mathsf{Q}}$. In addition to the abstract theory, we provide several tools for finding the numeraire in concrete cases. We discuss several nonparametric examples where we can indeed identify the numeraire and RIPr, despite not having a reference measure. We end with a more general optimality theory that goes beyond the ubiquitous logarithmic utility. We focus on certain power utilities, leading to reverse Rényi projections in place of the RIPr, which also always exist.
@@ 1 Introduction
Suppose we observe data from an unknown distribution on a measurable space $(\Omega,{\mathcal{F}})$. This paper is concerned with testing the composite null hypothesis $H_{0}$ that the data distribution belongs to a given set ${\mathcal{P}}$ of distributions on this space, against the point alternative $H_{1}$ that the data was generated by a given distribution ${\mathsf{Q}}$. Here ${\mathcal{P}}$ and ${\mathsf{Q}}$ are completely arbitrary, and our general theory will not require any further assumptions whatsoever.
We approach the testing problem by designing e-variables: a $[0,\infty]$-valued random variable $X$ is called an e-variable for ${\mathcal{P}}$ if ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Specifically, we will look for ‘optimal’ e-variables, in senses to be formalized later, that tend to be large under ${\mathsf{Q}}$.
E-variables measure evidence against the null: the larger its realized value, the stronger the evidence. One can also use e-variables to make hard accept/reject decisions. In particular, Markov’s inequality implies that we can reject $H_{0}$ at level $\alpha$ if an e-variable $X$ exceeds $1/\alpha$, since ${\mathsf{P}}(X\geq 1/\alpha)\leq\alpha$ for each ${\mathsf{P}}\in{\mathcal{P}}$. This paper will not be concerned with such hard decisions, but just in the measure of evidence $X$ itself, which requires no prespecification of any level $\alpha$.
@@
Our main result establishes that under absolutely no restrictions or conditions on ${\mathcal{P}}$ or ${\mathsf{Q}}$, there exists a special ${\mathsf{Q}}$-almost surely unique and strictly positive e-variable $X^{*}$ that we call the numeraire e-variable, or just the numeraire. It is characterized by the property that ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$. This in turn implies two other interpretable properties ${\mathbb{E}}_{\mathsf{Q}}[X^{*}/X]\geq 1$ and ${\mathbb{E}}_{\mathsf{Q}}[\log(X/X^{*})]\leq 0$, the second of which is called log-optimality. In fact, we show that $X^{*}$ is the numeraire if and only if it is log-optimal in this sense. The terminology derives from mathematical finance, where the numeraire portfolio is a central object analogous to the numeraire e-variable.
@@
Although no assumption on ${\mathcal{P}}$ or ${\mathsf{Q}}$ are needed in general, the theory simplifies if the alternative ${\mathsf{Q}}$ does not assign positive probability to any event that is null under every ${\mathsf{P}}\in{\mathcal{P}}$. In this case we say that ${\mathsf{Q}}$ is absolutely continuous with respect to ${\mathcal{P}}$, written ${\mathsf{Q}}\ll{\mathcal{P}}$. This natural generalization of absolute continuity for pairs of measures goes back to Halmos and Savage (1949) and is satisfied in most statistically relevant settings, while being significantly weaker than requiring a common dominating reference measure.
@@
The condition ${\mathsf{Q}}\ll{\mathcal{P}}$ has several consequences: (a) the numeraire is ${\mathsf{Q}}$-almost surely finite; (b) every e-variable is ${\mathsf{Q}}$-almost surely finite; and (c) the set of all e-variables for ${\mathcal{P}}$ is bounded in probability under ${\mathsf{Q}}$, meaning that $\lim_{a\to\infty}\sup_{X}{\mathsf{Q}}(X>a)=0$ where the supremum extends over all e-variables for ${\mathcal{P}}$. In fact, we show that the above properties are equivalent in that they imply each other, and they also imply that ${\mathsf{Q}}\ll{\mathcal{P}}$.
@@
Next, we associate to the numeraire a special sub-probability measure ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ defined via its density $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. This measure belongs to the bipolar of ${\mathcal{P}}$, denoted by ${\mathcal{P}}^{\circ\circ}$. The bipolar can be interpreted as the effective null hypothesis: it is the largest family of sub-probability measures that has the same set of e-variables as ${\mathcal{P}}$. These are the distributions against which e-variables for ${\mathcal{P}}$ do not provide any evidence.
@@
We show that ${\mathsf{P}}^{*}$ can be characterized in terms of properties normally associated with the reverse information projection (RIPr). Earlier work on the RIPr involve assumptions related to absolute continuity and/or finiteness of the minimum relative entropy between ${\mathsf{Q}}$ and ${\mathcal{P}}$, and in such settings the above ${\mathsf{P}}^{*}$ coincides with the RIPr. This includes, in particular, work by Grünwald et al. (2024) and Lardy et al. (2023), which in turn build on results in Li (1999). This leads us argue that ${\mathsf{P}}^{*}$ is the natural definition of the RIPr in the absence of any assumptions on ${\mathcal{P}}$ or ${\mathsf{Q}}$.
@@
One of our main contributions is to significantly relax the conditions required for a log-optimality theory of e-variables to be built, and for a RIPr to exist. Furthermore, in addition to the abstract theory we develop tools to aid in the computation of the numeraire and the RIPr in concrete examples. We establish several sufficient conditions to verify that an e-variable at hand is indeed the numeraire, and we provide several examples of nonparametric nulls ${\mathcal{P}}$ (such as bounded distributions, symmetric distributions and sub-Gaussian distributions) that illustrate how these sufficient conditions can be used. These examples are out of the reach of previous works due to their lack of a reference measure on which the aforementioned works centrally rely.
An important message of our work is that in concrete examples, the duality between the set of e-variables and the null hypothesis can be exploited in different ways depending on the nature of the null. For parametric nulls, the simplest approach is usually to first find the RIPr (and then the numeraire). For non-parametric nulls, it is often easier to first find the numeraire (and then the RIPr).
@@
We end the paper with a more general optimality theory that goes beyond the ubiquitous logarithmic utility, by studying other concave utilities and associated divergences. Without aiming for full generality, we focus on certain power utilities in place of the logarithm, which leads to reverse Rényi projections in place of the RIPr. These developments, again, take place under absolutely no conditions on ${\mathcal{P}}$ or ${\mathsf{Q}}$.
@@
Finally, we note that if ${\mathcal{P}}=\{{\mathsf{P}}\}$ is a singleton and ${\mathsf{Q}}\ll{\mathsf{P}}$, we obtain as a consequence that the numeraire is the likelihood ratio $d{\mathsf{Q}}/d{\mathsf{P}}$, which is well known to be log-optimal (Shafer, 2021). Thus, our work can be viewed as deriving certain generalized likelihood ratios for composite null hypotheses, and extending the corresponding optimality theory.
@@
Section 2 defines the numeraire, establishes existence and uniqueness under no assumptions, characterizes the absolute continuity property ${\mathsf{Q}}\ll{\mathcal{P}}$, and establishes various other useful facts such as log-optimality of the numeraire. Section 3 introduces the RIPr and characterizes it via various optimality properties dual to the properties and log-optimality of the numeraire. Section 4 establishes a number of tools for finding the numeraire and the RIPr in practice. These tools are then applied to specific examples in Section 5. Section 6 goes beyond log-optimality by considering certain power utilities and the associated reverse Rényi projections. Section 7 outlines implications for composite alternative hypotheses and draws comparison with universal inference. Section 8 concludes.
@@
Given the measurable space $(\Omega,{\mathcal{F}})$, we let $F_{+}$, $M_{+}$, and $M_{1}$ denote the set of all $[0,\infty]$-valued measurable functions, nonnegative measures, and probability measures on $\Omega$, respectively. Given a null hypothesis ${\mathcal{P}}$, which is an arbitrary nonempty subset of $M_{1}$, we write ${\mathcal{E}}$ for the associated set of e-variables, i.e., the family of all $X\in F_{+}$ that satisfy ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Infinite quantities play a key role in this paper. We generally rely on standard measure theoretic conventions such as $00\cdot x=$ 0 for $x\in[-\infty,\infty]$ and $x\cdot\infty=\infty$ for $x\in(0,\infty]$, as well as $\log(0)=-\infty$ and $\log(\infty)=\infty$. When evaluating ratios of $[0,\infty]$-valued random variables we often use the following convention, the last part of which is somewhat unusual but simplifies notation:
@@
Clarifying explanations are given in potentially ambiguous cases. Finally, we write $\operatorname{conv}(\cdots)$ for convex hull, that is, the set of (finite) convex combinations of the indicated objects.
@@ 2 The numeraire e-variable
A numeraire e-variable (or just numeraire for short) is a ${\mathsf{Q}}$-almost surely strictly positive e-variable $X^{*}$ such that ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every e-variable $X$.
@@
Numeraires are unique up to ${\mathsf{Q}}$-nullsets. Indeed, if $X^{*}_{1}$ and $X^{*}_{2}$ are numeraires, then the ratio $Y=X^{*}_{2}/X^{*}_{1}$ satisfies $1\leq 1/{\mathbb{E}}_{\mathsf{Q}}[Y]\leq{\mathbb{E}}_{\mathsf{Q}}[1/Y]\leq 1$ thanks to the numeraire property of $X^{*}_{1}$, Jensen’s inequality, and the numeraire property of $X^{*}_{2}$. Thus Jensen’s inequality holds with equality, so $Y$ is ${\mathsf{Q}}$-almost surely equal to a constant which must be one. It follows that $X^{*}_{1}$ and $X^{*}_{2}$ are ${\mathsf{Q}}$-almost surely equal. In view of this uniqueness, we often speak of ‘the’ numeraire.
@@
In the case of a simple null ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$ with ${\mathsf{Q}}$ absolutely continuous with respect to ${\mathsf{P}}_{0}$, the numeraire is just the likelihood ratio $X^{*}=d{\mathsf{Q}}/d{\mathsf{P}}_{0}$. Indeed, $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive, it is an e-variable, and for any other e-variable $X$ we have ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathsf{P}}_{0}}[X]\leq 1$. In particular settings, such as when a reference measure exists and certain additional assumptions are satisfied, we will show that the numeraire is a likelihood ratio involving the ‘reverse information projection’ (Grünwald et al., 2024; Lardy et al., 2023). Even without these restrictions, $X^{*}$ can still be interpreted as a likelihood ratio. One of our goals is to establish this in full generality. This is the subject of Section 3.
@@
If $X$ is an e-variable for ${\mathcal{P}}^{\prime}$, it is also an e-variable for ${\mathcal{P}}$ and one has ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ by assumption. This proves the lemma. ∎
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X}{X^{*}}\right]\leq 0$ for every e-variable $X$, (2.1)
where the left-hand side may be $\infty$- ∞. This follows directly from Jensen’s inequality and the numeraire property. The converse is also true, and we record the equivalence in the following proposition. The proof shows that the numeraire property is really the first-order condition for log-optimality. Note also that a numeraire $X^{*}$ is the ${\mathsf{Q}}$-almost surely unique log-optimal e-variable in the sense of (2.1) even if ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ happens to be infinite.
@@
Let $X^{*}$ be a ${\mathsf{Q}}$-almost surely strictly positive e-variable. Then $X^{*}$ is a numeraire if and only if it is log-optimal. In particular, a log-optimal e-variable is unique up to ${\mathsf{Q}}$-nullsets.
@@
The forward direction was argued above. To prove the converse we assume $X^{*}$ is log-optimal. For any e-variable $X$ and $t\in(0,1)$, $X(t)=(1-t)X^{*}+tX$ is an e-variable. Thus by log-optimality, ${\mathbb{E}}_{\mathsf{Q}}[t^{-1}\log(X(t)/X^{*})]\leq 0$. The expression inside the expectation equals $t^{-1}\log(1-t+tX/X^{*})$, which converges to $X/X^{*}-1$ as $t$ tends to zero and is bounded below by $t^{-1}\log(1-t)$, hence by $-2\log 2$ for $t\in(0,1/2)$. Fatou’s lemma thus yields ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}-1]\leq 0$, showing that $X^{*}$ is a numeraire. Finally, the uniqueness statement follows from the equivalence just proved together with uniqueness of the numeraire up to ${\mathsf{Q}}$-nullsets. ∎
The following theorem shows that a numeraire always exists without any assumptions whatsoever on ${\mathcal{P}}$ or ${\mathsf{Q}}$. The theorem is related to an important family of results in the mathematical finance literature, known collectively as the Fundamental Theorem of Asset Pricing, which are versions of the statement that a numeraire portfolio with finite wealth exists if and only if unbounded profits must come with unbounded downside risk. The following theorem contains the corresponding statement in the context of e-variables.
@@
    The numeraire is ${\mathsf{Q}}$-almost surely finite.
@@
    Every e-variable is ${\mathsf{Q}}$-almost surely finite.
@@
    ${\mathsf{Q}}$ is absolutely continuous with respect to ${\mathcal{P}}$.
@@
    The set of e-variables ${\mathcal{E}}$ is bounded in probability under ${\mathsf{Q}}$.
@@
The existence of a numeraire is Lemma 2.9 below, which relies on the preceding Lemmas 2.7 and 2.8. Here we focus on the equivalence of the four conditions. Clearly (ii) implies (i). To see that (i) implies (ii), let $X^{*}$ be a ${\mathsf{Q}}$-almost surely finite numeraire and note that the numeraire property implies that any e-variable $X$ must be ${\mathsf{Q}}$-almost surely finite too. Next, we claim that (ii) is equivalent to (iii). Suppose (iii) fails. Then there is an event $A$ with ${\mathsf{Q}}(A)>0$ and ${\mathsf{P}}(A)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$, in which case $X=\infty\bm{1}_{A}$ is an e-variable that is not ${\mathsf{Q}}$-almost surely finite. Thus (ii) fails also. The converse follows by noting that every e-variable is finite ${\mathsf{P}}$-almost surely for all ${\mathsf{P}}\in{\mathcal{P}}$. Further, (iv) implies (ii) because under the former condition we have for any e-variable $X$ that ${\mathsf{Q}}(X=\infty)=\lim_{a\to\infty}{\mathsf{Q}}(X>a)=0$. It remains to prove that (ii) implies (iv). This more involved implication is the content of Lemma 2.6 below. ∎
@@
The rest of this section contains the remaining parts of the proof of Theorem 2.5. This material can be skipped without compromising the understanding of Sections 3, 4, or 5.
@@
If every e-variable is ${\mathsf{Q}}$-almost surely finite then ${\mathcal{E}}$ is bounded in probability under ${\mathsf{Q}}$.